OpenStudy (anonymous):

g(x)=x^3-x^2 but is shifted right by 2 units. Write the equation for f(x).

OpenStudy (kc_kennylau):

f(x)=g(x-2)=(x-2)^3-(x-2)^2=you continue

OpenStudy (anonymous):

Thanks :)